OpenStudy (anonymous):

Part 1: Solve the inequality 12 + 4x less than or greater to 5x - 2 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

12 + 4x <= 5x - 2 4x <= 5x - 2-12 4x-5x <= -2-12 I'll let you finish up

OpenStudy (anonymous):

http://www.sketchtoy.com/29334463 Am I doing this right?

jimthompson5910 (jim_thompson5910):

-x <= -14 x >= 14 ... the sign flips because you're multiplying both sides by -1

OpenStudy (anonymous):

Okay, Thank you!

jimthompson5910 (jim_thompson5910):

yw, just gotta do the graph then you're done

OpenStudy (anonymous):

I think I can do the graph.

jimthompson5910 (jim_thompson5910):

ok great